# A Characterization for Tightness of the Sparse Moment-SOS Hierarchy

### v1

@@ 4.3. The case of finite sets
We prove the sparse Moment-SOS hierarchy of (3.2)-(3.3) is tight when the equality constraint of (4.6) defines a finite real variety.
@@
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.4). Assume (4.16) holds for all $i$ and let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
@@
Proof.
(i) Since each $V_{\mathbb{R}}(h_{i})$ is finite, there exists $N_{i}$ such that for all $\epsilon>0$,
$\in$ $\operatorname{Ideal}[h]_{spa,2N}+\operatorname{QM}[g]_{spa,2N}.$
$f_{\min}-m\epsilon\leq f_{N}^{spa}\leq f_{\min}.$
This forces $f_{k}^{spa}=f_{\min}$ for all $k\geq N$.
@@
The proof of Theorem 4.6 is given in Appendix A. The following result is straightforward by Theorems 4.5 and 4.6.
@@
Suppose the real variety $V_{\mathbb{R}}(h_{i})$ is a finite set for every $i\in[m]$ and $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP. Then, Assumption 4.1 holds and all conclusions of Theorem 4.5 hold.
@@ 5. The Schmüdgen type sparse relaxations
(5.1) $\left\{\begin{array}[]{rcl}f_{k}^{smg}\coloneqq&\max&\gamma\\ &\mbox{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{Pre}[g]_{spa,2k}.\end{array}\right.$
@@
(5.2) $\left\{\begin{array}[]{ccl}f_{k}^{smm}\coloneqq&\min&\langle f,y\rangle=\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ &\mbox{s.t.}&\mathscr{V}_{h_{i}}^{spa,2k}[y]=0\ (i\in[m]),\\ &&L_{g_{i,J}}^{spa,k}[y]\succeq 0\ (i\in[m],\,J\subseteq[s_{i}]),\\ &&y_{0}=1,\ y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$
We refer to (2.2) and (2.14) for the above notation. For a subset $J\subseteq[s_{i}]$, denote $g_{i,J}\coloneqq\prod_{j\in J}g_{i,j}.$ For the case $J=\emptyset$, $g_{i,\emptyset}=1$ and $L_{g_{i,\emptyset}}^{spa,k}[y]$ becomes the sparse moment matrix $M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]$. In the same way as for Theorem 3.1, we can get the following by Lemma 2.1.
@@
Now we consider the case that $K_{\Delta_{i}}$ is a finite set for every $i\in[m]$.
@@
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.4). Assume (5.5) holds for all $i$. Let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
@@
Proof.
(i) As we have shown in Theorem 4.4(i), the minimum value of (4.6) is 00. Since $K_{\Delta_{i}}$ is finite, there exists a degree $N_{i}$ such that for all $\epsilon>0$,
This is shown in the proof of Theorem 4.1 of [28]. Note that
$\in$ $\operatorname{Ideal}[h]_{spa,2N}+\operatorname{Pre}[g]_{spa,2N}.$
$f_{\min}-m\epsilon\leq f_{N}^{smg}\leq f_{\min}.$
This forces $f_{N}^{smg}=f_{\min}$, so $f_{k}^{smg}=f_{\min}$ for all $k\geq N$.
(5.8) $\left\{\begin{array}[]{cll}\min&\langle f_{i}+p_{i},y_{\Delta_{i}}\rangle\\ \mbox{s.t.}&\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]=0\,(i\in[m]),\\ &L_{g_{i,J}}^{\Delta_{i},k}[y_{\Delta_{i}}]\succeq 0\,(i\in[m],\,J\subseteq[s_{i}]),\\ &(y_{\Delta_{i}})_{0}=1,\ y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}^{\Delta_{i}}_{2k}}.\end{array}\right.$
Since each $K_{\Delta_{i}}$ is finite, by [30, Theorem 5.6.7], we know $y^{*}_{\Delta_{i}}$ satisfies the flat truncation (5.5) when $k$ is large enough. The remaining part follows from Theorem 3.3.
(iii) Note that $\mathcal{X}_{\Delta_{i}}\subseteq K_{\Delta_{i}}$, so $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ is also a finite set. Then, by the same proof as for Theorem 4.4(iii), we can get (5.6). ∎
We remark that when all $K_{\Delta_{i}}$ are finite sets and the RIP holds, then Schmüdgen type sparse Moment-SOS hierarchy of (5.1)-(5.2) is tight.
@@
Suppose all $K_{\Delta_{i}}$ are finite sets. If $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP, then Assumption 4.1 holds and all conclusions in Theorem 5.2 hold.
@@ 6. Some examples
(6.1) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}x_{2}(x_{1}^{2}+x_{2}-1)}_{f_{1}}+\underbrace{x_{2}^{2}x_{3}(x_{2}^{2}+x_{3}-1)}_{f_{2}}\\ \mbox{s.t.}&1-x_{1}^{2}-x_{2}^{2}\geq 0,\quad 1-x_{2}^{2}-x_{3}^{2}\geq 0,\end{array}\right.$
@@
(6.2) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}+4x_{1}x_{2}}_{f_{1}}+\underbrace{4x_{2}x_{3}-x_{3}^{2}}_{f_{2}}\\ \mbox{s.t.}&1-x_{1}^{2}-x_{2}^{2}\geq 0,\quad 1-x_{2}^{2}-x_{3}^{2}\geq 0,\end{array}\right.$
@@
(6.3) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}x_{2}x_{3}-x_{1}x_{2}}_{f_{1}}+\underbrace{x_{2}x_{3}x_{4}-x_{3}x_{4}}_{f_{2}}\\ \mbox{s.t.}&x_{1}^{2}=x_{1},\,x_{2}^{2}=x_{2},\,x_{3}^{2}=x_{3},\,x_{4}^{2}=x_{4},\\ &x_{1}+x_{2}+x_{3}\geq 1,x_{2}+x_{3}+x_{4}\geq 1,\end{array}\right.$
@@
(6.4) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=f_{1}(x_{1},x_{2})+f_{2}(x_{2},x_{3})+f_{3}(x_{1},x_{3})\\ \mbox{s.t.}&1-x_{1}^{4}-x_{2}^{4}\geq 0,\ 1-x_{2}^{4}-x_{3}^{4}\geq 0,\ 1-x^{4}_{1}-x_{3}^{4}\geq 0.\end{array}\right.$
@@
(6.5) $\left\{\begin{array}[]{cll}\displaystyle\min_{x\in\mathbb{R}^{n}}&\displaystyle\sum_{i=1}^{m}\big{[}x_{\Delta_{i}}^{T}Q_{i}x_{\Delta_{i}}+b_{i}^{T}x_{\Delta_{i}}\big{]}\\ \mbox{s.t.}&1-x_{\Delta_{i}}^{T}B_{i}x_{\Delta_{i}}-c_{i}^{T}x_{\Delta_{i}}\geq 0,\ i=1,\ldots,m.\end{array}\right.$
@@
(6.7) $\left\{\begin{array}[]{cll}\displaystyle\min_{x\in\mathbb{R}^{n}}&\displaystyle\sum_{i=1}^{m}\Big{[}b_{i}^{T}x_{\Delta_{i}}+x_{\Delta_{i}}^{T}Q_{i}x_{\Delta_{i}}+\big{(}{x_{\Delta_{i}}^{[2]}}\big{)}^{T}D_{i}x_{\Delta_{i}}^{[2]}\Big{]}\\ \mbox{s.t.}&1-c_{i}^{T}x_{\Delta_{i}}-x_{\Delta_{i}}^{T}B_{i}x_{\Delta_{i}}-\big{(}{x_{\Delta_{i}}^{[2]}}\big{)}^{T}H_{i}x_{\Delta_{i}}^{[2]}\geq 0,\ i=1,\ldots,m.\end{array}\right.$
@@
We set $m=n$ and each block $\Delta_{i}$ is given as in (6.6). All $b_{i},c_{i},Q_{i}$ and $B_{i}$ are randomly generated in the same way as in Example 6.5. Each matrix $D_{i}=R_{i}^{T}R_{i}$ is randomly generated with the entries of $R_{i}$ obeying the uniform distribution on $[0,1]$, and each matrix $H_{i}$ is generated in the same way. For these choices, the objective is a sum of SOS-convex polynomials and the constraining polynomials are SOS-concave. This can be shown as in [30, Example 7.1.4]. By Theorem 4.3, the sparse relaxation (3.2)-(3.3) is tight for all $k\geq 2$. We report the computational time (in seconds) for solving (6.7) by the sparse relaxation (3.2)-(3.3) and dense relaxation in [14] with $k=2$. In Table 2, we display the computational time by the sparse relaxation (3.2)-(3.3) on the left, and the time by the dense relaxation on the right, inside each parenthesis. When $n$ is $50$ or $100$, the dense relaxations cannot be solved since the computer is out of memory (oom).
@@
(6.8) $\left\{\begin{array}[]{cl}\displaystyle\min_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}+(x_{1}x_{2}-1)^{2}}_{f_{1}}+\underbrace{(x_{2}x_{3})^{2}+(x_{3}-1)^{2}}_{f_{2}}\\ \mbox{s.t.}&1-x_{1}^{2}\geq 0,\ 1-x_{2}^{2}\geq 0,\,1-x_{3}^{2}\geq 0.\end{array}\right.$

### v2

@@ 4.3. The case of finite sets
We discuss the case that each real variety $V_{\mathbb{R}}(h_{i})$ is finite. In the following, we prove the sparse Moment-SOS hierarchy of (3.2)-(3.3) is tight if the equality constraint of (4.6) defines a finite real variety for each $i$, under Assumption 4.1. For the dense case (i.e., $m=1$), similar results are shown in [30].
@@
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.5) and satisfies the RIP. Assume (4.12) holds for all $i$ and let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
@@
The proof of Theorem 4.5 is given in Subsection 7.4. Unlike the dense case, if we only assume that every real variety $V_{\mathbb{R}}(h_{i})$ is finite, then the sparse Moment-SOS hierarchy of (3.2)-(3.3) may not be tight (as shown in Example 3.4). The sparsity pattern there ($\Delta_{1}=\{1,2\}$, $\Delta_{2}=\{2,3\}$, $\Delta_{3}=\{1,3\}$) does not satisfy the RIP. However, if in addition, we assume the RIP, then Assumption 4.1 holds and the sparse Moment-SOS hierarchy (3.2)-(3.3) is tight, as shown below.
@@
The proof of Theorem 4.6 is given in Subsection 7.5. The following result is implied by Theorems 4.5 and 4.6.
@@
Suppose that $K\neq\emptyset$, each real variety $V_{\mathbb{R}}(h_{i})$ is a finite set, and $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP. Then, Assumption 4.1 holds and all conclusions of Theorem 4.5 hold.
@@ 5. The Schmüdgen type sparse relaxations
(5.1) $\left\{\begin{array}[]{rcl}f_{k}^{smg}\coloneqq&\max&\gamma\\ &\mathrm{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{Pre}[g]_{spa,2k}.\end{array}\right.$
@@
(5.2) $\left\{\begin{array}[]{ccl}f_{k}^{smm}\coloneqq&\min&\langle f,y\rangle=\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ &\mathrm{s.t.}&{V}_{h_{i}}^{spa,2k}[y]=0\ (i\in[m]),\\ &&L_{g_{i,J}}^{spa,k}[y]\succeq 0\ (i\in[m],\,J\subseteq[s_{i}]),\\ &&y_{0}=1,\ y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$
We refer to (2.2) and (2.14) for the above notation. For a subset $J\subseteq[s_{i}]$, denote $g_{i,J}\coloneqq\prod_{j\in J}g_{i,j}.$ For the case $J=\emptyset$, $g_{i,\emptyset}=1$ and $L_{g_{i,\emptyset}}^{spa,k}[y]$ becomes the sparse moment matrix $M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]$. Similar to Theorem 3.1, the theorem below follows from Lemma 2.1.
@@
For the dense case (i.e., $m=1$), if the feasible set is finite, the Schmüdgen type dense Moment-SOS hierarchy is tight; see [30, Theorem 4.1]. Interestingly, this result can be extended to the sparse case.
@@
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.5) and satisfies the RIP. Assume (5.5) holds for all $i$. Let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
@@
The proof of Theorem 5.2 is given in Subsection 7.6. The flat truncation (5.5) is sufficient for the dense Schmüdgen type moment relaxation (i.e., $m=1$) to be tight, while this condition (5.5) alone cannot guarantee tightness for the sparse case, because it may not imply (5.7).
To conclude this section, we show that the Schmüdgen type sparse Moment-SOS hierarchy of (5.1)-(5.2) is tight when all $K_{\Delta_{i}}$ are finite sets and the RIP holds. This is an interesting generalization of Theorem 4.1 in [30].
@@
Suppose all $K_{\Delta_{i}}$ are finite sets and $K\neq\emptyset$. If $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP, then Assumption 4.1 holds and all conclusions in Theorem 5.2 hold.
@@ 6. Some examples
(6.1) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}x_{2}(x_{1}^{2}+x_{2}-1)}_{f_{1}}+\underbrace{x_{2}^{2}x_{3}(x_{2}^{2}+x_{3}-1)}_{f_{2}}\\ \mathrm{s.t.}&1-x_{1}^{2}-x_{2}^{2}\geq 0,\quad 1-x_{2}^{2}-x_{3}^{2}\geq 0,\end{array}\right.$
@@
(6.2) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}+4x_{1}x_{2}}_{f_{1}}+\underbrace{4x_{2}x_{3}-x_{3}^{2}}_{f_{2}}\\ \mathrm{s.t.}&1-x_{1}^{2}-x_{2}^{2}\geq 0,\quad 1-x_{2}^{2}-x_{3}^{2}\geq 0,\end{array}\right.$
@@
(6.3) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}x_{2}x_{3}-x_{1}x_{2}}_{f_{1}}+\underbrace{x_{2}x_{3}x_{4}-x_{3}x_{4}}_{f_{2}}\\ \mathrm{s.t.}&x_{1}^{2}=x_{1},\,x_{2}^{2}=x_{2},\,x_{3}^{2}=x_{3},\,x_{4}^{2}=x_{4},\\ &x_{1}+x_{2}+x_{3}\geq 1,x_{2}+x_{3}+x_{4}\geq 1,\end{array}\right.$
@@
(6.4) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=f_{1}(x_{1},x_{2})+f_{2}(x_{2},x_{3})+f_{3}(x_{1},x_{3})\\ \mathrm{s.t.}&1-x_{1}^{4}-x_{2}^{4}\geq 0,\ 1-x_{2}^{4}-x_{3}^{4}\geq 0,\ 1-x^{4}_{1}-x_{3}^{4}\geq 0.\end{array}\right.$
@@
(6.5) $\left\{\begin{array}[]{cll}\displaystyle\min_{x\in\mathbb{R}^{n}}&\displaystyle\sum_{i=1}^{m}\big{[}x_{\Delta_{i}}^{T}Q_{i}x_{\Delta_{i}}+b_{i}^{T}x_{\Delta_{i}}\big{]}\\ \mathrm{s.t.}&1-x_{\Delta_{i}}^{T}B_{i}x_{\Delta_{i}}-c_{i}^{T}x_{\Delta_{i}}\geq 0,\ i=1,\ldots,m.\end{array}\right.$
@@
(6.7) $\left\{\begin{array}[]{cll}\displaystyle\min_{x\in\mathbb{R}^{n}}&\displaystyle\sum_{i=1}^{m}\Big{[}b_{i}^{T}x_{\Delta_{i}}+x_{\Delta_{i}}^{T}Q_{i}x_{\Delta_{i}}+\big{(}{x_{\Delta_{i}}^{[2]}}\big{)}^{T}D_{i}x_{\Delta_{i}}^{[2]}\Big{]}\\ \mathrm{s.t.}&1-c_{i}^{T}x_{\Delta_{i}}-x_{\Delta_{i}}^{T}B_{i}x_{\Delta_{i}}-\big{(}{x_{\Delta_{i}}^{[2]}}\big{)}^{T}H_{i}x_{\Delta_{i}}^{[2]}\geq 0,\ i=1,\ldots,m.\end{array}\right.$
@@
We set $m=n$ and each block $\Delta_{i}$ is given as in (6.6). All $b_{i},c_{i},Q_{i}$ and $B_{i}$ are randomly generated in the same way as in Example 6.5. Each matrix $D_{i}=R_{i}^{T}R_{i}$ is randomly generated with the entries of $R_{i}$ obeying the uniform distribution on $[0,1]$, and each matrix $H_{i}$ is generated in the same way. For these choices, the objective is a sum of SOS-convex polynomials and the constraining polynomials are SOS-concave. This can be shown as in [32, Example 7.1.4]. By Theorem 4.3, the sparse relaxation (3.2)-(3.3) is tight for all $k\geq 2$. We report the computational time (in seconds) for solving (6.7) by the sparse relaxation (3.2)-(3.3) and dense relaxation in [16] with $k=2$. In Table 2, we display the computational time by the sparse relaxation (3.2)-(3.3) on the left, and the time by the dense relaxation on the right, inside each parenthesis. When $n$ is $50$ or $100$, the dense relaxations cannot be solved since the computer is out of memory (oom).
@@
[…]